Consider two mutually exclusive new product launch projects that Nagano Golf is considering. Assume the discount rate for both products is 12 percent.

  

Project A: Nagano NP-30.
   Professional clubs that will take an initial investment of $700,000 at Time 0.
   Next five years (Years 1–5) of sales will generate a consistent cash flow of $300,000 per year.
   Introduction of new product at Year 6 will terminate further cash flows from this project.

  

Project B: Nagano NX-20.
   High-end amateur clubs that will take an initial investment of $850,000 at Time 0.
  

Cash flow at Year 1 is $250,000. In each subsequent year cash flow will grow at 10 percent per year.

   Introduction of new product at Year 6 will terminate further cash flows from this project.

  

Year NP-30 NX-20
0 –$ 700,000 –$ 850,000
1 300,000 250,000
2 300,000 275,000
3 300,000 302,500
4 300,000 332,750
5 300,000 366,025

  

Complete the following table: (Do not round intermediate calculations. Round your "PI" answers to 3 decimal places, e.g., 32.161, and other answers to 2 decimal places, e.g., 32.16. Enter your IRR answers as a percent.)

  

NP-30 NX-20
  Payback years years
  IRR % %
  PI
  NPV $ $
Answer #1
NP-30
Year Cash flow stream Cumulative cash flow
0 -700000 -700000
1 300000 -400000
2 300000 -100000
3 300000 200000
4 300000 500000
5 300000 800000
Payback period is the time by which undiscounted cashflow cover the intial investment outlay
this is happening between year 2 and 3
therefore by interpolation payback period = 2 + (0-(-100000))/(200000-(-100000))
2.33 Years
NX-20
Year Cash flow stream Cumulative cash flow
0 -850000 -850000
1 250000 -600000
2 275000 -325000
3 302500 -22500
4 332750 310250
5 366025 676275
Payback period is the time by which undiscounted cashflow cover the intial investment outlay
this is happening between year 3 and 4
therefore by interpolation payback period = 3 + (0-(-22500))/(310250-(-22500))
3.07 Years
NP-30
IRR is the rate at which NPV =0
IRR 0.322725326
Year 0 1 2 3 4 5
Cash flow stream -700000 300000 300000 300000 300000 300000
Discounting factor 1 1.322725 1.749602 2.314243 3.0611082 4.049005
Discounted cash flows project -700000 226804.5 171467.5 129632 98003.724 74092.27
NPV = Sum of discounted cash flows
NPV NP-30 = 0.000655393
Where
Discounting factor = (1 + IRR)^(Corresponding period in years)
Discounted Cashflow= Cash flow stream/discounting factor
IRR= 32.27%
NX-20
IRR is the rate at which NPV =0
IRR 0.21583157
Year 0 1 2 3 4 5
Cash flow stream -850000 250000 275000 302500 332750 366025
Discounting factor 1 1.215832 1.478246 1.797299 2.1852124 2.65685
Discounted cash flows project -850000 205620.6 186031.2 168308.1 152273.52 137766.5
NPV = Sum of discounted cash flows
NPV NX-20 = 1.2407E-07
Where
Discounting factor = (1 + IRR)^(Corresponding period in years)
Discounted Cashflow= Cash flow stream/discounting factor
IRR= 21.58%
NP-30
Discount rate 0.12
Year 0 1 2 3 4 5
Cash flow stream -700000 300000 300000 300000 300000 300000
Discounting factor 1 1.12 1.2544 1.404928 1.5735194 1.762342
Discounted cash flows project -700000 267857.1 239158.2 213534.1 190655.42 170228.1
NPV = Sum of discounted cash flows
NPV NP-30 = 381432.86
Where
Discounting factor = (1 + discount rate)^(Corresponding period in years)
Discounted Cashflow= Cash flow stream/discounting factor
NX-20
Discount rate 0.12
Year 0 1 2 3 4 5
Cash flow stream -850000 250000 275000 302500 332750 366025
Discounting factor 1 1.12 1.2544 1.404928 1.5735194 1.762342
Discounted cash flows project -850000 223214.3 219228.3 215313.5 211468.64 207692.4
NPV = Sum of discounted cash flows
NPV NX-20 = 226917.18
Where
Discounting factor = (1 + discount rate)^(Corresponding period in years)
Discounted Cashflow= Cash flow stream/discounting factor
NP-30
PI= (NPV+initial inv.)/initial inv.
=(381432.86+700000)/700000
1.54
NX-20
PI= (NPV+initial inv.)/initial inv.
=(226917.18+850000)/850000
1.27